### v1

# A Midsummer Meme’s Dream: Investigating Market Manipulations in the Meme Coin Ecosystem

Abstract.
From viral jokes to a billion-dollar phenomenon, meme coins have become one of the most popular segments in cryptocurrency markets. Unlike utility-focused crypto assets like Bitcoin or Ethereum, meme coins derive value primarily from community sentiment, making them vulnerable to manipulation. This study presents a cross-chain analysis of the meme coin ecosystem, examining 34,988 tokens across Ethereum, BNB Smart Chain, Solana, and Base. We characterize the tokenomics of meme coins and track their growth in a three-month longitudinal analysis. We discover that among high-return tokens (¿100%), an alarming 82.6% show evidence of extensive use of artificial growth strategies designed to create a misleading appearance of market interest. These include wash trading and a form of manipulation we define as Liquidity Pool-Based Price Inflation (LPI), where small strategic purchases trigger dramatic price increases. We also find evidence of schemes designed to profit at the expense of investors, such as pump and dumps and rug pulls. In particular, most of the tokens involved had previously experienced wash trading or LPI, indicating how initial manipulations often set the stage for later exploitation. These findings reveal that manipulations are widespread among high-performing meme coins and suggest that their dramatic gains are often likely driven by coordinated efforts rather than natural market dynamics.
Blockchain, Meme Coins, Market Manipulations
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacyccs: Security and privacy Economics of security and privacyccs: General and reference Measurementccs: Information systems Information retrieval
1. Introduction
The explosive rise of meme coins represents one of the most impressive trends in cryptocurrency markets. What started as a joke with Dogecoin in 2013 (Nani, 2022) has evolved into a serious market worth over \$117 billion by 2024 (CoinMarketCap, 2025c). Top meme coins like Dogecoin (\$53 billion), Shiba Inu (\$15 billion), and Pepe (\$8 billion) now rank among the most valuable cryptocurrencies, rivaling tokens created for functional use. Their appeal lies in the potential for extreme returns: Pepe has yielded gains of up to 400,000%, while Shiba Inu reached an incredible 80,000,000% at its peak (Coin, 2024). These extraordinary profits have attracted waves of investors hoping to find the next meme coin before its price explodes. This surge in interest has led to an unprecedented proliferation of meme coins, largely driven by platforms like pump.fun, which allow users to create tokens with minimal technical skills (Boaz Sobrado, 2025).
Despite their popularity, meme coins present unique vulnerabilities that distinguish them from other cryptocurrencies. Unlike utility tokens, meme coins rely exclusively on perceived interest to maintain their value (Nani, 2022), creating strong incentives for manipulations. Token creators and early investors can use artificial growth strategies to generate illusory activity. After attracting enough investors, these manipulators can execute profit extraction schemes, like pump and dumps (Xu and Livshits, 2019) or rug pulls (Zhou et al., 2024), to capitalize at the expense of investors. These manipulations make meme coin investments particularly risky, as trading activities may be fabricated to lure investors into projects designed to extract value from them.
This paper presents a comprehensive cross-chain analysis of the meme coin ecosystem, examining the economic characteristics of these tokens and the prevalence of market manipulations within high-performing coins. We collect data on 34,988 tokens from established cryptocurrency aggregators, simulating the information environment encountered by typical investors, and focus our research on four major blockchains: Ethereum, BNB Smart Chain (BSC), Solana, and Base. Our investigation begins with a detailed characterization of meme coin tokenomics across different blockchains, examining their price structures, supply strategies, and market capitalization patterns. We then perform a three-month longitudinal analysis of price growth, identifying 707 high-return tokens (¿100%). We uncover an alarming 82.6% of these high-performing meme coins exhibit clear evidence of market manipulation, including artificial growth strategies such as recurrent wash trading and what we define as Liquidity Pool-Based Price Inflation (LPI), an operation where strategic purchases with minimal capital investment trigger dramatic price increases. Furthermore, we find profit extraction schemes designed to capitalize on artificially generated interest, with pump and dumps and rug pulls executed after calculated waiting periods. Most significantly, we discover that 62.9% of tokens subjected to these exploitative operations had previously undergone wash trading or LPI operations, revealing how initial manipulations often serve as preliminary stages in broader exploitation strategies. We also find this pattern among meme coins delisted by established aggregators: tokens first use artificial growth tactics to meet listing requirements, then leverage their newfound visibility to execute exit strategies, maximizing profits before disappearing from the ecosystem. This connection between manipulation tactics is even more pronounced in delisted tokens, where 86.67% of those subjected to profit extraction had previously undergone artificial growth operations. These findings highlight the prevalence of manipulation strategies across high-return meme coins and demonstrate the need for improved investor protection mechanisms.
Our main contributions are the following:
  • Cross-chain Characterization: To the best of our knowledge, we provide the first comprehensive cross-chain analysis of meme coins, examining 34,988 tokens across Ethereum, BSC, Solana, and Base. We find blockchain-specific patterns in temporal evolution, price structures, supply strategies, and market capitalization distributions and perform a three-month longitudinal study of meme coin price growth.
  • Artificial Growth Detection: We detect strategies employed to create misleading appearances of market interest, including a form of manipulation we define as Liquidity Pool-Based Price Inflation (LPI). We find 82.6% of high-performing tokens exhibit concrete indicators of market manipulation, such as concentrated ownership structures, wash trading, or LPI activities, suggesting that the impressive gains seen in meme coins are often due to coordinated manipulation rather than natural market behavior.
  • Profit Extraction Manipulations Analysis: We identify techniques used to extract profits from tokens, including pump and dumps and rug pulls. We discover how these exploitation tactics in 62.9% of high-return involved tokens follow earlier growth manipulations (rising to 86.67% in delisted tokens), indicating how initial manipulations often set the stage for later exploitation.
2. Related Work
Memes and Meme Coins. Internet memes, defined by Shifman (Shifman, 2013) as “units of popular culture that are circulated, imitated, and transformed by individual Internet users,” have become significant cultural phenomena. Researchers studied meme propagation patterns (Zannettou et al., 2018), tracked their evolution using deep learning (Beskow et al., 2020), and analyzed how harmful content emerges through visual-textual combinations (Qu et al., 2023).
The intersection of meme culture and cryptocurrencies created meme coins, which differ from traditional cryptocurrencies through their community-driven nature, humor incorporation, and meme culture origins (Stencel, 2023; Nani, 2022). In particular, the U.S. Securities and Exchange Commission (SEC) defines meme coins as a “type of crypto asset inspired by internet memes, characters, current events, or trends for which the promoter seeks to attract an enthusiastic online community to purchase the meme coin and engage in its trading“ (Securities and Commission, 2025), a definition we adopt in this study. Kim et al. (Kim and Park, 2024) revealed how social media drive meme coin valuation, while Belcastro et al. (Belcastro et al., 2023) demonstrated that meme coin price movements require specialized prediction approaches. Long et al. introduced CoinCLIP (Long et al., 2024a), a framework leveraging visual-language models to classify meme coin viability, and released the Coin-Meme dataset (Long et al., 2024b) containing multimodal data from Solana-based meme coins.
Market Manipulations on Blockchains. The cryptocurrency ecosystem, characterized by its relative lack of regulatory oversight, has become susceptible to various forms of market manipulation (Gandal et al., 2018; Krafft et al., 2018; Cernera et al., 2023b; Daian et al., 2020; Cernera et al., 2024). These manipulative practices undermine market integrity and pose significant risks to investors.
Wash Trading is a manipulation that aims to create an illusion of market activity by generating artificial volume. Victor and Weintraud (Victor and Weintraud, 2021) found that up to 35% of ERC-20 token volume on certain exchanges consisted of wash trades. Similarly, Le Pennec et al. (Le Pennec et al., 2021) proposed machine learning approaches to identify wash trading activities across different blockchain networks. This manipulation technique has extended to non-fungible token (NFT) markets as well, with von Wachter et al. (von Wachter et al., 2022) and La Morgia et al. (La Morgia et al., 2023a) documenting significant wash trading activity in NFT marketplaces.
Pump and dump schemes represent one of the most prevalent manipulation strategies in cryptocurrency markets. These schemes involve inflating an asset’s price through misleading statements and coordinated buying (the ”pump”), followed by selling the overvalued assets (the ”dump”) once enough investors have been lured in (Kamps and Kleinberg, 2018). Xu and Livshits (Xu and Livshits, 2019) developed detection mechanisms for pump and dump operations on cryptocurrency exchanges, identifying over 300 instances of such manipulations. Complementing this work, other studies (Hamrick et al., 2019; La Morgia et al., 2023c, 2020) found that pump announcements typically generated average returns of 65% for organizers, mostly at the expense of unsuspecting investors. These works also examined the communication platforms used to coordinate such schemes, identifying Telegram and Discord as the main hubs, a finding also supported by Telegram-focused research (La Morgia et al., 2023b, [n. d.], 2025).
Lastly, rug pulls have become a harmful exit scam in decentralized finance (DeFi), where developers abandon projects after attracting investors. Zhou et al. (Zhou et al., 2024) analyzed 201 rug pulls and developed tools to detect associated risks, while Cernera et al. (Cernera et al., 2023b) identified ”1-day rug pulls” that generated \$240 million in profits, primarily on BNB Smart Chain.
This study investigates whether the extreme profitability observed in some high-performing meme coins arises from organic market dynamics or is driven by coordinated manipulations.
3. Methodology
To capture a comprehensive view of the meme coin ecosystem, we design a methodology consisting of three steps. We first identify tokens labeled as meme coins by established cryptocurrency platforms, then expand our dataset by collecting newly created tokens from additional sources, and finally apply classification techniques to distinguish meme coins from other cryptocurrency assets.
3.1. Data Collection
The meme coin ecosystem spans multiple blockchains, including well-known tokens and emerging projects. To capture this diversity, we adopt a blockchain-agnostic data collection strategy, sourcing information from mainstream crypto tracking platforms and specialized decentralized exchange (DEX) aggregators. This cross-chain approach enables us to observe how meme coin creators leverage different blockchains, track migration patterns, and identify platform preferences, uncovering cross-chain trends and insights that would be missed in a single-chain analysis.
3.1.1. Known Meme Coins
Our first approach focuses on established cryptocurrency aggregators that explicitly categorize tokens as meme coins, providing a foundation for verified meme coins. For this purpose, we collect data from three primary sources.
We start with CoinMarketCap (CoinMarketCap, 2025d) and CoinGecko (CoinGecko, 2025b), two leading cryptocurrency tracking platforms that maintain specialized meme token categories. CoinGecko defines meme coins as “Tokens with no intrinsic value, but with a strong community that fosters social interaction” (Coingecko, 2024), while CoinMarketCap describes them as “a type of cryptocurrency that originated from internet memes. They are often created as a joke or for the purpose of satire… Some of these coins have gained popularity due to their community-driven nature and the humorous themes they embody” (CoinMarketCap, 2025e). Both definitions align with the one provided by the SEC (Securities and Commission, 2025), which we adopt as the formal definition for our work. Using their free APIs (CoinMarketCap, 2025a; CoinGecko, 2025a), we collect data on 2,930 and 2,107 meme coins, respectively. Additionally, we utilize Gecko Terminal (CoinGecko, 2025c), which categorizes tokens by meme trends. This approach reflects the definition of meme coins as assets inspired by Internet trends and culture (Securities and Commission, 2025). We develop a custom scraper to collect data on 4,735 meme coins from pools tagged with relevant meme categories, including animal, dog, cat, inu, and pump.fun (a platform for meme coin creation).
By aggregating the results of these three services and removing duplicates, we obtain a total of 8,852 distinct meme coins.
3.1.2. Searching other Meme Coins
To expand our dataset, we implement a second approach focusing on newly created tokens. We develop scrapers for two DEX aggregators: DexScreener (Dexscreener, 2025) and CoinSniper (CoinSniper, 2025), which track real-time data on emerging tokens across decentralized exchanges. These platforms differ from CoinMarketCap and CoinGecko in their listing criteria. While the latter require stricter conditions, such as functional websites and listing on integrated exchanges (CoinGecko, 2025d; CoinMarketCap, 2025b), DEX aggregators implement more permissive policies: DexScreener automatically indexes tokens immediately upon the creation of the liquidity pool and the initial transaction activity (Screener, 2024), while CoinSniper requires only a basic application process, after which the projects appear immediately on their New Listings Page without further verification (CoinSniper.Net, 2025). Our scrapers collected 30,771 tokens from DexScreener and 44,611 from CoinSniper, resulting in 65,021 unique tokens.
Although these platforms are commonly used to monitor meme coins due to their focus on newly launched tokens on DEXs, they could list a wide range of token types, not just meme coins. This makes it necessary to define criteria for distinguishing meme coins from conventional cryptocurrencies, utility tokens, and other digital assets. To this end, we strictly adhere to the formal definition established by the SEC, which characterizes meme coins as ”crypto assets inspired by internet memes, characters, current events, or trends” (Securities and Commission, 2025). Our analysis revealed that these cultural and thematic associations are consistently embedded within token names, creating a distinctive linguistic pattern that can be systematically identified. Using this insight, we examine the names of known meme coins collected in Sec. 3.1.1 to identify common keywords. These keywords then served as markers for detecting additional meme coins within our larger dataset.
Name-Based Classification
To identify meme-related keywords, we apply Term Frequency–Inverse Document Frequency (TF-IDF) analysis to our verified meme token dataset. This technique highlights terms by assigning higher weights to words that appear frequently in meme coin names but are uncommon in general cryptocurrency naming (Ramos et al., 2003). First, in the preprocessing phase, we convert all token names to lowercase and remove special characters. Next, we tokenize these names into individual words (unigrams) and calculate TF-IDF scores for each word. To find the ideal number of keywords for our analysis, we use the elbow method, which identifies where additional keywords offer limited value in terms of relevance (Bholowalia and Kumar, 2014). As shown in Fig. 7 in the Appendix, the elbow analysis suggests 150 keywords as an optimal cutoff point. We then manually review these keywords, removing general cryptocurrency terms (like BSC, ETH, SOL) that appear because of cross-chain versions of the same token (fe.g., Dogecoin (SOL), Dogecoin (BSC)). This refinement process results in 126 meme-specific keywords. In particular, the final keyword set reveals several prominent patterns. Terms associated with successful meme coins show high frequency: dog appear in 7.24% of names, reflecting the influence of market leaders Dogecoin and Shiba-inu; cat in 6.92% of names; inu in 4.82%; pepe in 3.82%; and doge in 3.30%. We also find keywords related to contemporary trends, such as AI (4.96%) and Trump (2.39%).
We then apply this keyword-based classification to the dataset from Sec. 3.1.2, using the same preprocessing steps. We identify tokens as meme coins if they contain one or more of the 126 meme-specific keywords, resulting in the detection of 26,557 meme coins.
Pump.fun-Based Classification
Pump.fun, a Solana-based platform specialized in meme coins, has recently gained popularity as a launching platform for these tokens (Boaz Sobrado, 2025). It simplifies meme coin creation for non-technical users and features fair launch mechanisms through an initial bonding curve that raises token prices as purchases increase (pump.fun, 2025). In particular, a key characteristic of pump.fun tokens is their deterministic address format: all token addresses end with the string pump. This consistent naming convention provided a reliable method for identification. By leveraging this unique identifier, we discovered 4,140 meme coins, detecting a total of 28,975 different meme coins.
By combining the known meme coins collected in Sec. 3.1.1 with the newly identified tokens, we obtain a total of 35,220 meme coins.
3.1.3. Dataset Refinement
Manual inspection of high market cap or high-priced tokens in our dataset reveals some false positives, such as stablecoins or wrapped tokens. We apply a three-stage filtering process to eliminate misclassified tokens and retain only meme coins, minimizing false positives even at the cost of excluding some legitimate ones. This conservative approach prioritizes precision over recall, enhancing the reliability of our subsequent analyses.
First, we use CoinMarketCap’s API (CoinMarketCap, 2025a) to identify and remove stablecoins. By cross-referencing our dataset against CoinMarketCap’s list of 82 labeled stablecoins, we discard 4 tokens, reducing our dataset to 35,116 tokens.
In the second refinement stage, we focus on identifying non-meme tokens with significant market presence. Specifically, we examine tokens that meet either of two conditions: a price above \$0.80—since meme coins typically have very low prices (Nani, 2022), this threshold helps capture stablecoins that may be slightly depegged—or a market capitalization exceeding $10^{7}$, as the cumulative distribution function (CDF) in Fig. 8 (Appendix) shows that most tokens in our dataset fall below this value. This process identifies 525 tokens for detailed review. We then use CoinGecko’s API (CoinGecko, 2025a) to retrieve their associated categories and remove those labeled with non-meme classifications (e.g., Staking, Bridged). This step results in the removal of 46 additional tokens.
Finally, we apply a conservative string-matching filter to remove any remaining tokens with names containing terms typically associated with non-meme assets: specifically usd, wrapped, and staked. This step identifies and excludes 182 tokens, resulting in a final dataset of 34,988 meme coins. Tab. 1 provides an overview of the tokens collected from various sources and the filtering steps applied to refine the final dataset.
Table 1. Summary of meme coin data collection.
Source Tokens Collected
CoinMarketCap 2,930
CoinGecko 2,107
GeckoTerminal 4,735
Subtotal (Known Meme Coins) 8,852
CoinSniper 44,611
DexScreener 30,771
Subtotal (Other Tokens) 65,021
Name-Based Classification 26,557
Pump.fun Classification 4,140
Subtotal (Detected Meme Coins) 28,975
Total Meme Coins Before Filtering 35,220
Removed Stablecoins (CoinMarketCap) 4
Removed Non-Meme Tags (CoinGecko) 46
Removed Tokens (Wrapped/USD/Staked) 182
Final Meme Coin Dataset 34,988
[…]
This table summarizes the meme coins data collection and filtering process
3.2. Dataset Overview
We collected the data in mid-October 2024, capturing a comprehensive snapshot of the meme coin ecosystem. For each of the 34,988 tokens in our dataset, we gather technical details, including the token name, symbol, contract address, and blockchain. In particular, the contract address and blockchain serve as unique identifiers, enabling the tracking of tokens across different platforms.
3.2.1. Blockchain Distribution
The choice of blockchain significantly impacts the visibility and adoption of a meme coin. The 34,988 meme coins collected reveal clear trends in platform preference. BNB Smart Chain (BSC) hosts the largest share, with 16,405 tokens (46.88%), followed by Solana with 11,829 (33.80%), Ethereum with 3,814 (10.90%), and Base with 828 (2.36%).
Ethereum (Ethereum.org, 2025) was the first platform to support smart contracts, laying the foundation for decentralized applications. However, its high fees and scalability limitations led to the development of alternatives. One was the creation of new Layer 1 blockchains like BNB Smart Chain (BSC) (Bnbchain.org, 2024), which is EVM compatible and offers lower transaction costs. Another was the rise of Layer 2 solutions built on Ethereum itself, such as Base (Base.org, 2025), which improves scalability while remaining part of the Ethereum ecosystem. In contrast, Solana (Foundation, 2025b) is not EVM compatible and takes a different approach, prioritizing low fees and high transaction throughput through its unique proof-of-history consensus mechanism. Together, these four blockchains account for 32,876 meme coins, representing 93.96% of our dataset. Given their dominance, our subsequent analysis focuses on tokens deployed on these platforms.
3.2.2. Addresses Validation
As the next step, we perform a validation process to identify and discard invalid addresses and non-deployed tokens (e.g., placeholders such as 0x000comingsoon or upcoming). For EVM-compatible blockchains (BSC, Ethereum, and Base), we first validate address format: they must be 42 characters long, start with ”0x,” and contain only hexadecimal characters. This step identifies 791 invalid addresses (739 in BSC, 51 in Ethereum, 1 in Base). Next, for addresses with the correct format, we use the explorers’ APIs (Etherscan (Etherscan, 2025) for Ethereum, BscScan (BscScan, 2025) for BSC, and BaseScan (BaseScan, 2025) for Base) to check for their creation transactions. We find 254 addresses, 211 in BSC and 43 in Ethereum, that lacked creation transactions and, thereby, are not deployed contracts. For Solana tokens, we use Solana’s RPC API (Foundation, 2025a) to verify valid mint addresses, identifying 20 invalid addresses.
Tab. 2 summarizes the validation process and its results. After removing the 1,065 invalid and non-deployed addresses, the final dataset contains 31,811 tokens (96.76%): 15,455 on BSC, 11,809 on Solana, 3,720 on Ethereum, and 827 on Base.
Table 2. Summary of token address validation process
Blockchain Invalid Not Deployed Removed Remaining
BSC 739 211 950 15,455
Ethereum 51 43 94 3,720
Solana - 20 20 11,809
Base 1 0 1 827
Total 791 254 1,065 31,811
3.2.3. Temporal Evolution of Meme Coin Creation.
To analyze the creation patterns of the meme coins over time, we collect their contract deployment timestamps. For EVM-compatible blockchains, we use their explorers’ APIs (Etherscan, BscSca, and BaseScan). Concerning Solana, we leverage the API offered by Alchemy (Alchemy Insights, 2025), a blockchain development service that provides tools for building decentralized applications.
Fig. 1 shows the temporal distribution of meme coin creation across different blockchains from 2017 to 2024. The data highlights how platform preferences for meme coin deployment have shifted over time, from Ethereum’s early lead to BSC’s rise in popularity and, more recently, to the growing adoption of Solana and Base. During the initial period (2017–2020), meme coin activity remained low, with fewer than 10 new tokens created per month. This scenario changed significantly due to several key events.
Late 2020 Bull Run. The cryptocurrency bull market, marked by Bitcoin’s rise from \$10,000 to over \$60,000 (Martin Young, 2021), coincided with two major developments: the launch of SHIBA INU and the introduction of Binance Smart Chain (BSC), that significantly boosted meme coin creation thanks to its low transaction fees (Coin98, 2023).
Pepe Coin & BONK. In April 2023, the launch of Pepe coin on Ethereum challenged BSC’s dominance. This was followed by the surge of BONK in October 2023, driven by the introduction of single-sided staking pools (Compass, 2024).
Pump.fun. The peak in meme coin creation came in the summer of 2024, especially in June and July, when Solana emerged as the leading platform. This surge is largely due to the rapid rise in popularity of pump.fun (Boaz Sobrado, 2025).
Figure 1. Creation date of meme coins over time and their distribution on different blockchains.
[…]
A plot showing the creation date of meme coins over time and their distribution on different blockchains.
4. Tokenomics and Price Growth
In this section, we group meme coins by their creation dates and analyze their economic characteristics across different blockchains to uncover distinctive patterns within the meme coin ecosystem. We then examine their price movements over three months to understand how their value changes over time.
4.1. Tokenomic Anatomy Across Blockchains
[…]
A figure containing three CDF plots comparing meme coins across Ethereum, BSC, Solana, and Base. The first plot shows the distribution of price, the second shows the distribution of circulating supplies, and the third shows the distribution of market cap.
(a) Price
(b) Circulating Supply
(c) Market Cap
Figure 2. Distribution of the prices (2(a)), circulating supplies (2(b)), and market cap (2(c)) of the meme coins on Ethereum, BSC, Solana, and Base grouped according to their creation date.
Of the 31,811 meme coins collected, we successfully obtained economic information for 20,692 meme coins (65.05%) in US dollars. The remaining 11,119 tokens for which we cannot retrieve additional data have been collected on DexScreener, Coinsniper, and CoinGecko. These platforms derive token price, circulating supply, and market capitalization from liquidity pools where tokens are traded. However, if a token lacks a liquidity pool at the time of data collection, its price and market cap cannot be determined. This explains why no such information is available for these 11,119 tokens. After removing four tokens with a price of zero (all on BSC), the final dataset consists of 20,688 meme coins: 10,468 on Solana, 7,493 on BSC, 2,107 on Ethereum, and 620 on Base.
We group the meme coins in our dataset according to when they were created to examine how their age correlates with their economic performance and characteristics. Specifically, we group tokens into quarterly periods for 2024 ($Q1_{2024}$, $Q2_{2024}$, $Q3_{2024}$, and $Q4_{2024}$), with an additional category for tokens created before 2024 (pre-2024). Fig. 2 illustrates how economic metrics vary across these periods for each blockchain platform.
4.1.1. Price Structures
The price analysis reveals an ecosystem characterized by tokens with very low prices, alongside a few high-priced outliers. The overall distribution shows a median price of \$$4\times 10^{-6}$ with approximately 85% of tokens priced below \$0.00001. This pricing strategy creates a psychological illusion of accessibility: investors can acquire billions of tokens for minimal capital, fostering the narrative that these assets could replicate the trajectories of Dogecoin or Shiba Inu (Nani, 2022). Moreover, we identify 14 meme coins with prices exceeding \$100, each employing one of two distinct price strategies. The first involves artificial scarcity mechanisms using regular token burning, as seen in tokens like Ponzio The Cat (valued at approximately \$20 billion) and HayCoin (valued at \$250,000). These tokens have been flagged as scams by security platforms such as Token Sniffer (Sniffer, 2025) and show systematic supply reduction designed to inflate prices, rewarding early investors at the expense of those who enter later. The second approach involves the use of single-token supply models, as seen in tokens such as Potato (valued at \$55,000). This extreme scarcity is leveraged to justify their unusually high valuations, despite limited utility or demand.
Concerning the temporal analysis of the price, Fig. 2(a)) reveals three distinct blockchain-specific pricing patterns.
Higher Prices for Newer Tokens on Ethereum and BSC: These platforms exhibit sequential price increases across time periods, suggesting a market preference for newly launched tokens. Pre-2024 tokens show median prices of $6\times 10^{-8}$ (Ethereum) and $5\times 10^{-10}$ (BSC), while $Q4_{2024}$ tokens reach $4\times 10^{-5}$ (Ethereum) and $1\times 10^{-4}$ (BSC), representing three to six orders of magnitude increase.
Cyclical pricing on Base: Tokens on Base demonstrate non-monotonic price progression, starting with relatively high values in early periods ($2\times 10^{-5}$ for pre-2024), declining through mid-2024 (to $2\times 10^{-6}$ in $Q3_{2024}$), and recovering to $1\times 10^{-5}$ by $Q4_{2024}$.
Stable pricing on Solana: Conversely, Solana demonstrates remarkable price consistency across all creation periods, maintaining a median of approximately $8\times 10^{-6}$ regardless of token age.
4.1.2. Token Supply Strategies
An important element of the tokenomics of a crypto token is its circulating supply, the total number of tokens currently available on the market (not locked or burned). While established cryptocurrencies like Bitcoin (21 million coins) and Ethereum (about 120 million) emphasize supply scarcity, meme coins often implement hyperinflated supply models: 43.29% of meme coins in our dataset present circulating supplies exceeding 1 billion tokens, with 30.68% exceeding 100 billion tokens. Even at the 25th percentile, meme coin supplies reach 973 million tokens. This extreme supply inflation serves a psychological purpose: it allows promoters to highlight the appeal of owning billions of tokens, distracting from their actual low value. As shown in Fig. 2(b), supply strategies vary significantly across blockchains.
Hyperinflated supplies on Ethereum and BSC: These platforms host tokens presenting high variability with median supplies often exceeding $10^{11}$.
Moderate supplies on Base: Base tokens typically maintain supplies in the $10^{9}$-$10^{11}$ range, with significant but less extreme variation compared to Ethereum and BSC.
Standardized supplies on Solana: Solana exhibits consistent token supplies concentrated around $10^{9}$, due to the standardized creation processes implemented by pump.fun, which applies fixed supply parameters.
4.1.3. Market Capitalization
Market capitalization, calculated by multiplying the circulating supply of a token by its current price, is a key indicator of the market value of a project. Among the meme coins we analyzed, the median market cap is just \$4,100, highlighting their highly speculative nature. Interestingly, a few meme coin projects report extremely high market capitalizations. While some well-known tokens, such as Dogecoin (\$53 billion), Shiba Inu (\$15 billion), and Pepe (\$8 billion), have substantial and legitimate market value, these are rare exceptions in a space often marked by value manipulation. For example, catgirl a claims an implausible market cap of \$151 billion despite having almost no liquidity, suggesting how unreliable market capitalization can be as a standalone metric for meme coins. Upon investigating the 47 meme coins reporting market caps above \$1 billion, we detect significant evidence of market manipulation. More than half have liquidity under \$1,000, and 88.1% show liquidity below 0.1% of their reported market cap. This discrepancy points to a common tactic: inflating token prices through small, strategic trades in low-liquidity environments, a form of manipulation detailed in Sec. 5.3.
Regarding the cross-chain analysis, Fig. 2(c) suggests consistent temporal patterns in market cap across all blockchains: newer tokens (those created in $Q4_{2024}$) present dramatically higher market caps compared to older ones. Specifically, Base shows the highest median market capitalization for recent tokens at \$882,507, followed by Ethereum at \$443,688 and Solana at \$214,981. BSC tokens remain significantly lower at \$12,223, approximately one order of magnitude below their counterparts on other blockchains.
This strong preference for newly launched tokens is more consistent across blockchains than pricing patterns, suggesting that token novelty may be a stronger driver of investor capital allocation than the underlying blockchain infrastructure.
4.1.4. Cross-Platform Economic Patterns
Our analysis of price, supply, and market capitalization reveals that each blockchain has its own economic pattern. Ethereum meme coins show higher prices for newer tokens, extremely large token supplies, and high market caps. BSC presents a similar pattern to Ethereum but with key differences. While it shares hyperinflated token supplies and rising price trends for newly created tokens, BSC meme coins maintain consistently lower market caps than all other platforms. Conversely, Solana demonstrates remarkably consistent pricing regardless of token age, standardized supply structures, and moderate market caps, reflecting its more structured and possibly regulated approach to meme coin creation. Instead, Base exhibits characteristics of an emerging ecosystem with cyclical price volatility rather than consistent trends, moderate supply metrics, and dramatic variations in market capitalization.
4.2. Three-Month Price Performance Evaluation
To study the price evolution of meme coins, we tracked their prices over three months following our initial data collection until mid-January 2025.
4.2.1. Data Availability and Retrieval
Of the 20,688 meme coins we analyzed in Sec.4.1, we successfully obtained updated price data for 20,327 tokens (98.25% of the original sample), including 10,438 on Solana, 7,355 on BSC, 1,922 on Ethereum, and 612 on Base.
Of the 361 tokens with missing price data, we originally collected these from multiple sources: 12 from Gecko Terminal, 350 from DexScreener, and 21 from Coingecko (with some overlap between platforms). As discussed in Sec. 4.1.1, platforms such as Gecko Terminal and DexScreener require tokens to maintain active liquidity pools to determine prices. Thus, the absence of price data for these tokens likely indicates their liquidity pools were removed during our study period. Meanwhile, Coingecko may delist tokens that fail to meet its listing criteria (CoinGecko, 2025d), particularly those with poor liquidity or absent development teams. The blockchain with the highest percentage of tokens lacking updated price data is Ethereum with 185 tokens (8.78%), followed by BSC with 138 (1.84%), Base with 8 (1.29%), and Solana with 30 (0.29%).
4.2.2. Price Change Distribution
Fig. 3 presents the cumulative distribution functions (CDFs) of meme coin price changes across different blockchains. For clarity, tokens with returns exceeding 100% are capped at 100% in the visualization. Specifically, 6,938 tokens (34.13%) experience price decreases, 6,462 tokens (31.79%) maintain stable prices, and 6,927 tokens (34.07%) see price increases.
Figure 3. CDFs of meme coin price changes by blockchain platform over three months.
[…]
A plot showing the CDFs of the price changes of meme coins grouped by different blockchains (Ethereum, BSC, Solana, and Base).
Negative Return Tokens. The 6,938 tokens that decreased in value lost an average of 38.42%, with a median loss of 27.97%. Substantial losses are common across the dataset: 946 tokens (13.63%) lost more than 80% of their value, while 2,299 tokens (33.13%) lost more than 50% of their value. In particular, performance varies substantially across blockchains. Ethereum shows the largest median price decrease at 42.13%, followed by BSC and Base with median decreases of 38.19% and 36.95%, respectively, while Solana tokens demonstrated greater resilience with a median decrease of only 26%. However, when analyzing the proportion of tokens with price losses within each blockchain, Solana presents the highest rate (50.43% of its tokens), followed by Base (46.41%), Ethereum (40.95%), and BSC with the lowest rate (8.21%).
Inactive Tokens. Interestingly, 6,462 tokens (31.79% of the sample) maintained exactly the same price over the three-month period. This price stability appears to result from complete inactivity rather than market equilibrium, as all of these tokens recorded zero trading volume during our observation period. Unlike the 361 tokens for which price updates were unavailable, services like DexScreener continue recording the prices of these inactive coins as they maintain open liquidity pools. Regarding the distribution of inactive tokens across blockchains, the majority are on BSC, with 5,277 tokens (71.75% of BSC tokens with updated price). Ethereum follows with 609 inactive tokens (31.69%), then Solana with 523 tokens (5.01%), and Base with 53 tokens (8.66%).
Positive Return Tokens. Among the 6,927 tokens that increased in value, extreme outliers skewed the mean increase to $10^{35}$%, while the median is only 21.46%. The distribution of positive returns shows that 707 tokens (10.21%) gained more than 100%, and 171 tokens (2.47%) achieved returns exceeding 1,000%. Across blockchains, Ethereum tokens present the highest median price increase at 29.33%, while tokens on other chains averaged approximately 20% growth. The proportion of tokens with positive price performance is highest on Base and Solana (44% each), followed by Ethereum (25%), and lowest on BSC (19.65%).
If we exclude inactive tokens (those with a stable price), Fig. 3 suggests that investing in meme coins could result in positive growth in about 50% of cases, sometimes yielding significant profits. However, this analysis overlooks two key factors. First, it does not account for the 361 tokens for which updated price data was unavailable—many of which likely lost all their value. Second, as discussed in Sec. 4.1.3, meme coins with high returns may be subject to market manipulations that artificially inflate their perceived value. To explore this further, we examine the 707 tokens that exhibited exceptionally high returns (¿100%) after three months.
5. Artificial Growth Strategies
This section analyzes high-performing meme tokens: First, anomalous activity patterns; then, evidence of manipulations designed to distort the market. In particular, in order to analyze potential manipulations, we collected OHLCV (Open, High, Low, Close, Volume) data with hourly granularity from our initial data collection in mid-October 2024 through mid-January 2025 using the Gecko Terminal APIs (GeckoTerminal, 2025).
5.1. Market Anomalies in High-Return Tokens
To better understand the behavior of the 707 meme coins that exhibit returns exceeding 100%, we investigate potential anomalies in their market characteristics. To support this study, we collected additional data from Gecko Terminal and DexScreener, including ownership concentration metrics and security scores from monitoring services such as GoPlus (GoPlus, 2025), De.Fi (De.Fi, 2025), and Soul Scanner (Scanner, 2024).
5.1.1. Ownership Concentration Anomalies
Ownership concentration occurs when a small number of wallets control most of the token supply. This concentration poses significant risks for investors, as major holders can manipulate prices through coordinated trading activities and other manipulative practices, effectively controlling the token’s value at the expense of smaller investors who often lack visibility into these ownership structures. We detect severe ownership anomalies across many high-return meme coins, with various techniques employed sometimes to concentrate token control while creating an illusion of distributed ownership.
Concentrated Holdings. We identify 385 tokens (54.46%) with an alarming concentration of ownership among a small number of holders. The top 10 holders controlled an average of 77.85% of these tokens, with a median of 87.01%. Following standards used by established blockchain security platforms such as GoPlus (GoPlus, 2025) and CertiK (CertiK, 2025), we identify problematic concentration when the top holders control more than 30% of a token’s supply. This is the case for 359 tokens (50.78% of high-return tokens).
Bundle Buy Concealment. A strategy developers can use to obscure their token holdings is leveraging bundle buys (e.g., JITO bundles (Explorer, 2025)), an approach that allows them to group multiple transactions together and execute them in a specific order. This enables the creation of bundles that first add liquidity (depositing tokens into a liquidity pool) and then immediately swap tokens, ensuring trades are executed right after liquidity is added. By distributing the swapped tokens across multiple wallets, they can mask true ownership and create the appearance of distributed holdings. We observe 133 tokens (18.81%) utilizing bundle buys, with an average of 15.70% of their supply involved. As for the concentrated holdings, we consider a bundle buy an anomaly when it exceeded 30% of a token’s supply, which occurred in 20 cases (2.83%).
Fresh Address Distribution. Another concealment technique is the use of multiple newly created addresses. This tactic allows token creators or top holders to give the illusion of decentralized ownership while maintaining control for coordinated manipulation. We detect eight projects where a significant portion of tokens (more than 30%) are held in fresh addresses, wallets with no prior blockchain history. On average, these addresses control 40.25% of the token supply, with a median of 38%.
Strategic Airdrop Distribution. Airdrops involve distributing tokens for free to multiple addresses and are commonly used for legitimate promotional purposes. However, they can also serve as a concentration tactic. By strategically distributing tokens to multiple controlled wallets, token creators can create the illusion of decentralized ownership while maintaining effective control. The average percentage of tokens distributed via airdrops is 63.43%, with a median of 79.36%. In particular, 24 tokens (3.40%) have more than 30% of their total supply allocated by airdrops.
5.1.2. Honeypots
Unlike ownership concentration anomalies which focus on token distribution, honeypots represent a trading activity anomaly. These fraudulent tokens appear legitimate but are designed with malicious code that prevents investors from selling after purchase, effectively trapping their funds. This deceptive structure allows creators to collect investment capital while ensuring investors cannot exit their positions, regardless of how much the token’s value drops. We find 28 meme coins flagged as honeypots by Go Plus Token Security (GoPlus, 2025), representing a significant risk to unwary investors seeking high returns in the meme coin market.
5.1.3. Anomaly Prevalence
Tab. 3 provides a summary of the statistics of the anomalies detected. Of the 707 high-performing tokens, 412 meme coins (58.27%) exhibit at least one anomaly indicator. Among these, 27 tokens show two anomalies: 26 combine honeypot behavior with ownership concentration, while one displays both airdrop activity and bundle buys. The most common issue is ownership concentration, where the top 10 holders own over 30% of the supply, occurring in 87.14% of the detected anomalies. Concerning the other indicators, although they are less frequent, they still help in flagging suspicious tokens that might otherwise escape detection.
Table 3. Anomaly statistics for high-performing meme coins.
Anomaly Type # tokens Percentage in anomalies
Top Holders 359 (50.78%) 87.14%
Honeypots 28 (3.96%) 6.80%
Airdrop 24 (3.39%) 5.83%
Bundle Buys 20 (2.83%) 4.85%
Fresh Addresses 8 (1.13) 1.94%
Total tokens 412 (58.27%) 100%
[…]
The table shows the anomaly statistics for meme tokens with high returns, including the number and percentage of tokens affected by each type of anomaly.
The prevalence of anomalies in high-performing tokens raises serious doubt on the authenticity of their market success. To establish more concrete proof of artificial market activity, the following two subsections examine manipulation techniques that directly impact token valuation and visibility.
5.2. Wash Trading Analysis and Detection
Wash trading is a deceptive practice in which the same entities buy and sell tokens back and forth to create an illusion of market activity. This manipulation inflates the trading volume without real ownership changes or market risk, creating a false perception of popularity and demand (Le Pennec et al., 2021). Meme coins are especially vulnerable to this strategy, as their success often depends on social sentiment and perceived popularity (Nani, 2022).
5.2.1. Initial Detection Methodology
To detect potential wash trading, we identify suspicious trading patterns where volume increased dramatically (¿500% compared to the previous day) while the price remained nearly unchanged (absolute price change ¡5%). This pattern contrasts with natural market behavior, where significant volume surges typically drive notable price movements. We select the 500% volume threshold to focus only on extreme trading activity that deviates from normal market patterns, while the 5% price stability criterion helps distinguish these suspicious activities from legitimate high-volume trading that would typically cause significant price changes. Applying this method to the OHLCV data, we detect 3,264 potential wash trading instances involving 377 tokens.
5.2.2. Advanced Detection Techniques
To strengthen our wash trading detection and provide more concrete evidence, we use the Dune Analytics API (Dune, 2025) to collect detailed transaction data. This enables us to identify the specific traders (makers) involved in suspicious transactions on the days flagged by our initial methodology. We apply three different techniques to detect wash trading activities.
Zero-Risk Position. Wash trading, by definition, involves zero-risk market manipulation where traders buy and sell without taking actual market risk. To detect this strong evidence of wash trading, we look for makers who bought and sold almost identical amounts of the same token on the same day. To account for slippage and transaction fees, we allow for a small 2% difference between buy and sell volumes, consistent with methodologies used in previous works (La Morgia et al., 2023a). This approach detects 958 manipulations (29.35% of suspected cases), 271 involved tokens (71.88% of suspected tokens), and 1,777 unique makers.
Circular Volume. A strong indicator of wash trading is when nearly all trading volume comes from the same set of addresses acting as both buyers and sellers. Such a high concentration of volume among circular traders strongly suggests coordinated manipulation rather than genuine activity. Thus, we analyze instances where at least 99% of daily trading volume is generated solely by makers who both bought and sold the token on the same day. This technique identifies 219 (6.71%) suspicious activities involving 128 (33.95%) tokens and 670 unique makers engaged in circular trading.
Persistent Manipulator Network. The presence of persistent manipulator networks indicates organized and sustained efforts to manipulate the market. To uncover potential collusion networks, we examine whether the same makers participated in multiple wash trading events for the same token, focusing specifically on identifying persistent manipulators who appeared in every detected wash trading day for a given token. This technique detects 37 manipulations involving 14 tokens and 55 unique makers.
Tab. 4 summarizes the results of the wash trading detection analyses. By combining the three detection approaches, we identify 1,014 events presenting strong evidence of wash trading activities (31.07% of suspected instances) and 282 tokens affected (74.8% of suspected tokens). The blockchain distribution of the manipulated tokens reveals that most are on Solana, with 181 tokens, followed by BSC with 52, Ethereum with 28, and Base with 21 tokens.
Table 4. Wash trading detection results by methodology
Analysis Method Events Detected Tokens Affected
Zero-Risk Position 958 (94.48%) 271 (96.10%)
Circular Volume 219 (21.60%) 128 (45.39%)
Persistent Manipulators 37 (3.65%) 14 (4.96%)
Total Detected Cases 1,014 282
The overlap of the detection techniques reveals that most wash trading activities (75.44%) are uniquely identified by zero-risk position, while only smaller proportions are exclusively detected through circular volume (4.04%) or persistent manipulator (1.38%). In particular, some activities exhibit multiple patterns: 17.46% are flagged by both zero-risk and circular volume techniques, while only 0.69% display all three wash trading indicators simultaneously. At the token level, the zero-risk position detects 94.48% of the meme coins suspected of wash trading, suggesting this approach forms the foundation of most wash trading operations.
5.2.3. Wash Trading Characteristics
After detecting instances of wash trading, we examine their frequency, volume characteristics, and temporal distribution to understand how manipulators create artificial market interest.
Recurrent Manipulations. Affected tokens experience an average of 3.59 wash trading operations, with a median of 3 and a maximum of 22, a result consistent across all blockchains. This pattern suggests recurring manipulation rather than isolated activities. Multiple wash trading operations on the same token likely serve to sustain the illusion of legitimacy and create a perception of growing interest over time.
Volume Generated. Wash trading operations create dramatic spikes in trading volume compared to the previous day, showing a median increase of 1,759% and an average increase distorted by extreme outliers reaching $10^{8}$%. During wash trading events, the average absolute trading volume is \$6,174, with a median of \$830, presenting platforms-specific differences. BSC shows the lowest median volume (\$366), followed by Base (\$516) and Solana (\$928), while Ethereum records the highest median volume (\$1673). This pattern likely reflects Ethereum’s higher transaction fees, which make small-volume wash trades economically impractical compared to other platforms with lower gas costs.
Temporal Patterns and Lifecycle Analysis.
Figure 4. Temporal distribution of wash trading activity per token, relative to its creation date, across different blockchains.
[…]
The figure shows a scatterplot of wash trading activities occurring on each meme coin since its creation date, with different colors representing different blockchains.
Fig. 4 illustrates the timeline of wash trading activity for meme coins. Most tokens experience their first manipulation within a median of 85 days, with some targeted earlier: 6.03% (17) within their first week and 24.47% (69) within their first month. Specifically, Solana tokens are targeted earlier, with a median of just 62 days after creation, compared to medians of at least 101 days on other blockchains. Moreover, these wash trading campaigns often continue for extended periods, averaging 87 days with a median of 64 days, indicating prolonged and sustained manipulation. Only 39.36% (111) of manipulated tokens experience activity for 30 days or less. For tokens subjected to multiple manipulations, the average time between consecutive operations is 33 days, with a median of 16 days. Specifically, 19.92% (202) of wash trading events occur within at most one week, and 49.41% (501) occur within at most one month. This pattern holds across blockchains, except for BSC, which has a slightly shorter median interval of 9 days.
Few Wash Traders. Wash trading activities typically involve very few participants, with an average of 2.84 actors. In particular, 513 operations (50.59%) were performed by a single actor. This concentrated manipulation pattern suggests coordinated market manipulation carried out by a small group of entities or even individuals, a trend also observed in previous studies on wash trading (La Morgia et al., 2023a).
These repeated wash trading operations performed by a few entities on the same tokens over extended periods suggest a strategic purpose: creating trading volume to maintain the appearance of legitimate market activity. Specifically, wash trading can serve two strategic purposes. First, it helps tokens reach the minimum trading volume required for listing on established cryptocurrency aggregators (CoinGecko, 2025d; CoinMarketCap, 2025b), ensuring visibility to potential investors. Second, it increases the ranking of a meme coin on platforms such as DexScreener (Dexscreener, 2025), which sorts tokens by trading volume. This increased visibility attracts unsuspecting investors, who perceive the artificial volume as evidence of genuine market interest.
5.3. Liquidity Pool-Based Price Inflation (LPI)
While wash trading distorts market perception through artificial volume, we identify another form of manipulation that directly targets token price rather than activity metrics. We define this technique as Liquidity Pool-Based Price Inflation (LPI), a manipulation that exploits the mechanics of liquidity pools on decentralized exchanges. LPI leverages the constant product formula ($x\times y=k$) used by liquidity pools to compute token prices. Indeed, in low-liquidity pools, even small trades can trigger disproportionate price movements. Manipulators exploit this mechanism by executing strategic purchases that dramatically inflate a token’s apparent market value with minimal investment, creating a deceptive impression of rapid appreciation that attracts unsuspecting investors.
To illustrate this manipulation, consider a simplified scenario: A liquidity pool contains 1 ETH and 1000 MEME tokens, making each MEME worth 0.001 ETH. When a manipulator purchases 500 MEME tokens for 0.5 ETH, the pool composition changes to 500 MEME and 1.5 ETH. Following the constant product formula, the token’s price now is 0.003 ETH, a 200% increase achieved through just a single transaction.
5.3.1. LPI Detection
To detect LPI activities, we develop a conservative two-phase methodology. The first phase focuses on identifying statistically anomalous trading patterns by applying two main criteria: (1) a significant price increase exceeding 100% within a 24-hour period, which may indicate artificial price inflation; and (2) either minimal growth in trading volume compared to the previous day ($\leq 20\%$ increase) or an absolute trading volume below \$1,000, both of which distinguish synthetic price movements from legitimate market demand. This approach isolates the defining characteristic of LPI operations, namely, abrupt price increases that lack corresponding trading activity. When applied to our OHLCV dataset, this screening process identifies 871 events involving 149 tokens.
The second phase implements a more stringent verification step to minimize false positives. Using the Dune Analytics API, we retrieve granular transaction-level data for each flagged event and introduce two additional constraints. First, buy volume must account for at least 90% of the total daily trading volume, indicating sustained buying pressure rather than balanced market activity. Second, the number of unique trading entities must not exceed 10, implying coordinated or strategic behavior rather than organic participation by a broad user base. This methodology detects 72 price manipulation events affecting 40 distinct meme coins, 12 on Solana, 11 on BSC, 10 on Ethereum, and 7 on Base.
5.3.2. LPI Analysis
Next, we examine LPI events focusing on their financial impact, actor participation, and the targeted tokens.
Figure 5. Relationship between price increase percentage and buy volume for LPI activities across different blockchains.
[…]
The figure shows a scatter plot showing the relationship between price increase percentage (x-axis) and buy volume (y-axis) for LPI activities across different blockchains.
High Price Impact with Minimal Investment. Fig. 5 shows the relationship between buy volume and price impact across the detected events revealing that manipulators achieved substantial price increases with minimal investment. Indeed, the median buy volume is merely \$$\$54$, with only $43.06\%$ (31) of events requiring more than \$$\$100$ in capital. Ethereum stands apart from other blockchains with its \$560 median, while others show median values below \$5. Despite these modest volumes, the effects on the prices are dramatic: $50\%$ of manipulations generate price increases exceeding $500\%$, while $36.11\%$ produce gains surpassing $1,000\%$. In particular, the most extreme price distortions (exceeding $10,000\%$) occur with the smallest capital inputs, highlighting the vulnerability of low liquidity pools. The manipulation effectiveness varied significantly across blockchains: Base demonstrated the highest median price increase at $4\times 10^{8}\%$, followed by Solana ($661\%$), BSC ($504\%$), and Ethereum ($214\%$).
Few Actors Manipulation. The analysis of participant patterns further confirms the coordinated nature of these manipulations. On average, each LPI operation involves just 1.96 participants, with 66.67% of events carried out by a single actor. This trend is consistent across all blockchains studied, highlighting how vulnerable low-liquidity pools are to price manipulation by individual actors.
Targets Are Old Tokens. Unlike other manipulation techniques that frequently target newly launched tokens (Cernera et al., 2023a, b), LPI occurs predominantly at the more mature stages of the lifecycle of a token. Only 2 meme coins experience manipulation within their first month after launch, and only 9 tokens are manipulated within their first 90 days. The median time to manipulation is 209 days after token creation, suggesting that manipulators specifically seek out older tokens that never gained significant traction. These tokens present ideal targets as they appear more legitimate due to their longer existence while still maintaining the low liquidity pools necessary for effective price manipulation. This temporal pattern is particularly different on Solana, where tokens are targeted much earlier, with a median of just 115 days, compared to tokens on other blockchains, which have a median of over 163 days.
These results reveal that LPI is one of the most capital-efficient market manipulation strategies. With a relatively small investment, a single actor can drive significant artificial price increases, producing misleading price charts that attract unwary investors looking for high-growth opportunities.
5.4. Legitimate vs. Manipulated Growth
The analysis of high-performing tokens shows that the vast majority display signs of market manipulation or conditions that facilitate it. By combining results from wash trading (282 tokens), LPI strategies (40), and strong anomaly indicators (412), we identify 584 tokens (82.6% of all high-performing tokens) that exhibit at least one form of suspicious activity. In particular, 117 tokens (20.03%) show wash trading and strong anomaly indicators, with 114 tokens exhibiting extreme ownership concentration due to the top 10 accounts holding an average of 73.09% of the supply. Ownership concentration is even more pronounced in cases related to LPI, where of the 40 involved tokens, 32 show clear signs of concentrated ownership. This includes 27 tokens where the top 10 holders control an average of 80.62% of the supply, and two cases where all the supply is acquired through bundle buy transactions, likely to obscure actual ownership. Furthermore, we observe 11 tokens involved in both wash trading and LPI activities. This combination of tactics suggests coordinated efforts to inflate trading volume and price, creating a convincing illusion of strong market demand and growth.
This analysis leaves only 123 tokens that show no signs of market manipulation or suspicious ownership. In terms of blockchain distribution, these successful tokens are primarily on Solana (82 tokens), followed by BSC (22), Ethereum (12), and Base (7). When normalized by the total number of tokens on each platform for which we have price data (see Section 4.2.1), Base has the highest success rate (with 1.14%), followed by Solana (0.79%), Ethereum (0.62%), and BSC with the lowest success rate (0.30%).
6. Profit Extraction Manipulations
Following our discovery that 82.6% of high-performing meme coins exhibited manipulative practices (Sec. 5.4), this section explores whether such artificial growth is part of a calculated strategy to extract profits at the expense of investors. To track this progression from manipulation to extraction, we extended our analysis with OHLCV data from mid-January through mid-March 2025. This additional window allows us to observe whether tokens that experienced price increases are later involved in exit strategies.
6.1. Strategic Exit via Rug Pulls
Rug pulls represent a severe form of exit scam where developers abandon their project after generating significant initial investment, typically by removing all liquidity from trading pools (Zhou et al., 2024). To detect rug pulls among high-performing tokens, we used two strict conservative criteria: (1) a single-day price drop of over 99%, signaling that the token has effectively lost all value, a decline too extreme to be explained by normal market fluctuations, and (2) a sustained drop in trading volume of more than 99% in the following days, indicating that liquidity was removed and investors could no longer trade the token. These conditions confirm that the token has been abandoned and has not experienced a temporary market dip.
Applying this approach to our updated OHLCV data, we detect two tokens subjected to apparent rug pulls: one on Solana and the other on Ethereum. The timing of these events reveals calculated planning, with developers executing rug pulls 248 and 451 days after token creation, respectively. This suggests a strategy of establishing credibility over an extended period before executing the exit. Moreover, both tokens have previously undergone other market manipulations to generate artificial interest, one involved in wash trading and the other in price pumping, indicating a progressive manipulation strategy culminating in the final exit scam.
6.2. Orchestrated Pump and Dump Operations
Pump and dumps are another common manipulation in which the price of a token is artificially inflated through coordinated buying, followed by rapid selling that crashes the price and leaves late investors with substantial losses (Xu and Livshits, 2019). Our methodology for detecting pump and dump operations is the following:
1. Price peaks detection: We identify price peaks demonstrating overbought conditions with a Relative Strength Index exceeding 80, indicating excessive buying momentum (Ţăran-Moroşan, 2011). For each detected peak, we examine the preceding 24-hour window to locate the local minimum price point as the potential manipulation starting point.
2. Pump phase: We classify a pump phase when three conditions are simultaneously met: (a) price increase exceeding 50% from starting point to peak, (b) trading volume surge exceeding 500% during this period, and (c) complete price ascent occurring within a maximum 24-hour timeframe.
3. Dump phase: We then detect the subsequent dump phase through a sharp price decline exceeding 30% following the peak, with post-dump average trading volume falling below 50% of the volume observed during the pump phase, indicating rapidly waning interest after the orchestrated collapse.
To minimize false positives, we implemented more conservative thresholds than previous research by Kamps et al. (Kamps and Kleinberg, 2018), who employed a 10% price increase threshold with a 400% volume surge requirement. This approach detects 91 pump and dump operations affecting 60 different tokens, 54 on Solana, 3 on BSC, 2 on Base, and 1 on Ethereum.
Figure 6. Relationship between price increase during pump and decrease during dump across different blockchains.
[…]
Scatterplot showing the magnitude of price increases during pump phases (x-axis, logarithmic scale) and subsequent price decreases during dump phases (y-axis) for each pump and dump operation, color-coded by blockchain.
As shown in Fig. 6, the pump phases exhibit significant variability in magnitude. The median price increase is 143.95%, with 63.74% of operations (58) featuring increases of at least 100% and 7 operations (7.69%) with increases exceeding 1,000%. The subsequent dump phases show correspondingly severe price collapses, with a median decrease of 55.82%. In 59.34% of cases (54), tokens lost at least 50% of their value, while 17.58% (16) experienced severe dumps exceeding 80% price decline. Lastly, concerning the time elapsed between the creation date of meme coins and their first pump and dump, we have that, on average, passes 206 days, with a median of 146 days and a minimum of 49 days.
The 60 affected tokens reveal a concerning pattern: 61.67% (37 tokens) were previously involved in artificial growth strategies, 34 in wash trading, and 3 LPI. If we also consider the rug pulled tokens, this ratio increases to 62.9%. This finding strongly suggests that manipulations examined in Sec. 5 often function as preliminary steps in a broader scheme designed to attract investors before profit extraction through pump and dumps or rug pulls.
6.3. Case Study: Delisted Tokens
The removal (delisting) of tokens from major aggregators like CoinMarketCap and CoinGecko can signal either project abandonment or suspicious market behavior. To better understand manipulation tactics across this ecosystem lifecycle, we analyze tokens that were once listed, having met the platforms’ stringent listing criteria, but were later removed. These tokens represent a particularly interesting case study, as they gained the credibility and visibility associated with major platform listings before their removal. Our investigation focuses on two key questions: whether these tokens employed artificial growth techniques to satisfy initial listing criteria, and whether they subsequently exploited their enhanced visibility through market manipulations.
We examine 124 meme coins delisted from CoinMarketCap and CoinGecko between October 2024 and March 2025 and report the results in Tab. 5. Among these, 84 tokens (67.74%) showed signs of artificial growth manipulation, with 70 tokens exhibiting wash trading behavior, while 27 tokens employed LPI tactics. Additionally, 30 tokens (24.19%) show evidence of profit extraction manipulation after listing: 23 tokens present pump and dump patterns, while 7 are involved in rug pulls. Of these, 26 tokens (86.67%) had previously employed artificial growth tactics. Among the pump and dump tokens, 14 showed evidence of wash trading and 5 LPI. Likewise, all rug-pulled tokens showed signs of prior manipulation, with 5 involving wash trading and 6 LPI (4 of which exhibited both).
Table 5. Manipulation strategies in delisted meme coins
Manipulation Category # Tokens Percentage
Wash Trading 70 56.45%
Liquidity Pool-Based Price Inflation (LPI) 27 21.77%
Tokens with Artificial Growth 84 67.74%
Pump and Dump 23 18.55%
Rug Pull 7 5.65%
Tokens with Profit Extraction 30 24.19%
These results highlight a recurring manipulation strategy among delisted meme coins. Tokens first achieve legitimacy by leveraging artificial growth manipulations to meet listing thresholds. After gaining the visibility that comes with the listing, they execute exit strategies, revealing a coordinated effort to maximize gains before disappearing from the ecosystem.
7. Discussion and Ethical Considerations
Our study reveals that market manipulations are widespread among high-performing meme coins, posing significant risks to unsuspecting investors. DEX aggregators are failing to protect and inform users adequately. While services like DexScreener and Gecko Terminal offer security metrics, they place critical warnings at the bottom of pages, where they are easily missed by users making quick investment decisions. Moreover, technical metrics like token concentration percentages are difficult for average users to understand. Even more concerning is that these platforms do not provide alerts about specific manipulation tactics such as wash trading or LPI, which are often used to simulate fake interest and inflate token values. To improve user protection, these platforms should redesign their interfaces to highlight risk indicators prominently at the top of token pages. These indicators should be clear, simple, and designed to help everyday users understand when a token displays signs of manipulation. This issue becomes even more pressing when considering that we observe recurring manipulation strategies on tokens previously listed on established aggregators such as CoinMarketCap and CoinGecko. These platforms, viewed as reliable due to their stricter listing requirements, inadvertently lend credibility to manipulated tokens. To better safeguard users, these aggregators must enhance their verification processes by introducing checks that detect artificial growth strategies, tactics we have identified as early steps in broader profit extraction schemes.
Concerning ethical considerations, all data examined in this work is publicly available on blockchains, and the addresses are pseudo-anonymous. Throughout our investigation, we made no attempts to correlate addresses with external information that could compromise user privacy. Consequently, according to our IRB’s policy, we did not need explicit authorization to perform our study.
8. Limitations
The nature of our cross-chain analysis, encompassing four different blockchains, made direct on-chain data extraction computationally prohibitive. This limitation is particularly pronounced for high-throughput blockchains like Solana, where complete transaction history analysis would require substantial infrastructure investment. To overcome these constraints while maintaining analytical rigor, we leveraged third-party APIs: Gecko Terminal for ecosystem-wide OHLCV data, and public nodes and Dune Analytics for transaction-level examination of high-performing tokens.
To prioritize precision over recall, our market manipulation detections employ deliberately conservative thresholds selected to identify clear cases where trading patterns significantly deviate from natural market behavior. For instance, a 500% volume surge with minimal price movement represents an anomaly difficult to attribute to legitimate trading. This conservative approach introduces a detection bias toward more extreme manipulation cases, potentially missing subtler distortions. However, this methodological choice significantly increases our confidence in the validity of the detected manipulations and ensures that our manipulation rates represent a reliable lower bound rather than speculative estimates.
9. Conclusion and Future Work
Meme coins have rapidly expanded into the cryptocurrency market. However, their community-driven nature makes them highly susceptible to manipulation. This study presents the first large-scale cross-chain analysis of meme coins, covering 34,988 tokens across four major blockchains. We found that 82.6% of high-performing meme coins exhibited clear signs of artificial growth strategies, including wash trading and LPI schemes, and concentrated token ownership. This result suggests that the dramatic gains often associated with meme coins are more likely the result of coordinated manipulation than organic market dynamics. We also detected pump and dumps and strategic rug pulls, with the majority of the tokens involved having previously experienced artificial growth tactics, indicating that early manipulations often serve as precursors to larger exploitation strategies. Finally, we also documented this manipulation pattern in tokens delisted from established aggregators, discovering that 67.74% showed signs of artificial growth manipulation before delisting, and 24.19% exhibited profit extraction tactics after gaining visibility through listing.
These findings underscore the urgent need for stronger investor protection measures in the meme coin ecosystem. Future work should focus on developing real-time detection systems capable of identifying artificial growth tactics such as wash trading and LPI operations, as well as tracking how these evolve into profit extraction schemes. Finally, to gain a deeper understanding of manipulation strategies, it would be valuable to examine the off-chain promotional activities of projects involved in market manipulation. This includes analyzing their websites and social media to understand their strategies and study how online engagement is coordinated with the timing of manipulative activities.
Appendix A Appendix
Figure 7. TF-IDF scores of keywords associated with meme coin names.
[…]
A plot showing the TF-IDF scores of keywords extracted from meme coins’ names.
Figure 8. CDF of the market capitalization of meme coins collected.
[…]
A plot showing the CDF of the market capitalization of meme coins collected.